OpenStudy (anonymous):

functions f(x)= sq.40x and g(x)= sq. 10x a) (f+g)= b)(f-g)= c)(fg)= d)(f/g)=

OpenStudy (anonymous):

think f(x) as f and g(x) as g a. (f+g)= sq.40x+sq10x then solve and you get 3sq10x b/c 2 is a perfect sq you can pull it out of 4. Then you will have 2sq10x+sq10x=3sq10x So, b. sq40x-sq10x same as in part a 2sq10x-sq10x=sq10x c. sq40x*10x=20x d. sq40x/sr10x=2